OpenStudy (cccmpt2):

How do you factor 2x^2-8?

jimthompson5910 (jim_thompson5910):

first factor out the GCF 2 2x^2 - 8 = 2(x^2 - 4)

jimthompson5910 (jim_thompson5910):

what's your next step?

OpenStudy (cccmpt2):

i have no idea! i don't know how to factor these at all!

jimthompson5910 (jim_thompson5910):

are you familiar with the difference of squares rule?

OpenStudy (cccmpt2):

no, i am studying for a placement test. so, most of the problems i have never learned anything about.

jimthompson5910 (jim_thompson5910):

it's the idea that a^2 - b^2 factors to (a-b)(a+b)

jimthompson5910 (jim_thompson5910):

so x^2 - 4 can be written as x^2 - 2^2

jimthompson5910 (jim_thompson5910):

then you can see that a = x and b = 2

jimthompson5910 (jim_thompson5910):

so what does that mean in this case?

OpenStudy (cccmpt2):

sorry, i'm lost

jimthompson5910 (jim_thompson5910):

x^2 - 4 can be written as x^2 - 2^2 does that make sense?

OpenStudy (cccmpt2):

i found it in my husband's old college algebra book, so i am gonna look it over now. thank you for your help ;)

jimthompson5910 (jim_thompson5910):

alright, anyways you would factor x^2 - 4 to get (x-2)(x+2) using the difference of squares rule above

jimthompson5910 (jim_thompson5910):

so 2(x^2 - 4) turns into 2(x-2)(x+2)

OpenStudy (cccmpt2):

I think I got it! 50-2x^2=2(5-x)(5+x) :)

OpenStudy (cccmpt2):

that was actually really easy! thank you for your help jim_thompson5910 :)